A 20.0-kg particle has a deBroglie wavelength of 10.0 cm. How fast is the particle moving?

A. 1.30 x 10-34 m/s

B. 2.64 x 10-34 m/s

C. 3.31 x 10-34 m/s

D. 3.98 x 10-34 m/s

E. 4.40 x 10-34 m/s


1
Expert's answer
2021-06-01T16:24:44-0400
"mv=\frac{h}{\lambda}\\(20)v=\frac{6.63\cdot10^{-34}}{0.1}\\v=3.31\cdot10^{-34}\frac{m}{s}"